A juggler throws next ball when previous ball is at maximum height .If he throws two ball per second then maximum height attained by each ball is
Ans- 1.25m

Grade:12th pass

1 Answers

Arun
5 years ago

If juggler throws n balls each second then-

Final Velocity = v = 0

Initial Velocity = u 

Time = 1/n seconds

0 = u - g*1/n

g/n = u

Now applying third equation of motion,

v2 = u2 - 2gH

0 = g/n * g / n - 2gH

-(g/n)2 = - 2gH

g2/n2 = 2gH

g/n2 = 2H

H = g/2n2

Now

n = 2

Then

H = g/8 = 10/8

H = 1.25 m